OpenStudy (anonymous):

What does 8y+4 equal?

OpenStudy (loser66):

with this question , how can I answer? it's ambiguous, friend. which class are you?

OpenStudy (littlenugget):

i think maybe she meant solve for y?

OpenStudy (anonymous):

idk but i am looking for the value of y here is the full problem If RS= 8Y+4,ST =4y+8, and RT= 15y-9, find the value of y

OpenStudy (loser66):

oh yeah, it makes sense now, however, I still need the pic.

OpenStudy (anonymous):

12y

OpenStudy (littlenugget):

 Is this what it looks like?

OpenStudy (anonymous):

here

OpenStudy (littlenugget):

ohhhh

OpenStudy (loser66):

hahahaah.... help him, please @littlenugget

OpenStudy (littlenugget):

 hopefully this drawing can help :) so RT is the whole line RS ST are just parts of the line. So.... RS+ST=RT they gave you the equations for each, so you would just plug them in: (8y+4)+(4y+8)=15y-9 So you are just going to solve for y.